OpenStudy (anonymous):

Find the equation of the tangent line to the following curve at point (0,pi) sin(x+y)=2x I know that to find the equation you will have to take the derivative which will give you the slope which you can then use along with the points to find the equation. I think that the derivative of the equation is cos(x+y)=2 so cos(x+y)-2=0 would you have to use the chain rule in the center?

OpenStudy (anonymous):

d/dx [sin(x+y)] = d/dx [ 2x] (1+dy/dx) * cos(x+y) = 2 helps ?

OpenStudy (anonymous):

you didnt differentiate it correctly

OpenStudy (anonymous):

So wait, is it now (2/cos(x+y))-1=dy/dx and then that is the slope?

OpenStudy (anonymous):

yes.. it means y'(x) = (2/cos(x+y))-1 so plug in here the x value of the point and youll have the slope

OpenStudy (anonymous):

do you understand how i took the derivative ?

OpenStudy (anonymous):

Okay so then m is 2/((cospi)-1) right?

OpenStudy (anonymous):

no m is (2/cos(pi))-1

OpenStudy (anonymous):

so then would the equation be 2/cospi-1x +pi?

OpenStudy (anonymous):

first simplify (2/cos(pi))-1 you should know what is this number

OpenStudy (anonymous):

OKay so that would be -1?

OpenStudy (anonymous):

what is cos(pi) ?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

so again what is (2/cos(pi))-1 ?

OpenStudy (anonymous):

\[\frac{ 2 }{ \cos(\pi) } - 1\]

OpenStudy (anonymous):

so then it would be -2-1 so -3?

OpenStudy (anonymous):

yes -3

OpenStudy (anonymous):

sorry for not answering

OpenStudy (anonymous):

That is alright Thank you for your help!

OpenStudy (anonymous):

yw:)